There are 24 families living in the Willbrook Farms Development. Of these families, 12 prepared their own federal income taxes for last year, 4 had their taxes prepared by a local professional, and the remaining 8 by H&R Block.

  1. What is the probability of selecting two families, neither of which had their taxes prepared by H&R Block? (Round your answer to 4 decimal places.)
Answer #1

Answer:

Given,

To give the probability of selecting two families, neither of which had their taxes prepared by H&R Block

Consider,

P(A and B) = P(A)*P(B|A)

P(A) = 20C1 / 24C1

= 20/24

P(B|A) = 19C1 / 23C1

= 19/23

P(A and B) = (20/24) / (19/23)

= 20/24 * 23/19

= 1